OpenStudy (soahelp):

What is the solution to the equation 1.6m − 4.8 = −1.6m?

OpenStudy (anonymous):

1.6m − 4.8 = −1.6m 3.2m = 4.8 Can you complete the solution from this point?

OpenStudy (soahelp):

would the answer be 0.7

OpenStudy (kendricklamar2014):

To complete the solution, we do:\[\left(\begin{matrix}3.2m \\3.2\end{matrix}\right) = \left(\begin{matrix}4.8 \\ 3.2\end{matrix}\right)\]

OpenStudy (soahelp):

so it would be 1.5=m

OpenStudy (kendricklamar2014):

yes the answer would be m = 1.5

OpenStudy (soahelp):

thank you

OpenStudy (kendricklamar2014):

No problem